OpenStudy (anonymous):

2^x + 3^y = 5, 2^(x+3) - 3^(y-2) = 23 solve these simultaneously??

OpenStudy (anonymous):

Suppose u can just proceed as normal to eliminate one variable. Say multiply the first one by 2^3 to get 2^(x+3) and then subtract to get y

OpenStudy (zarkon):

\[\alpha+\beta=5\] \[\alpha\cdot2^3-\beta\cdot3^{-2}=23\] solve this linear system of equations